What is the length of AB in the graph....I can give the coordinates..
B is -1,-3
A is 4,5
Answers are:
1) 53
2) 53 sqrt 2
3) 2 sqrt 53
4) sqrt 106
Thank you!

Answer by venugopalramana(3286)

What is the length of AB in the graph....I can give the coordinates..
B is -1,-3
A is 4,5
AB=SQUARE ROOT OF{(5-(-3))^2+(4-(-1))^2}=SQRT(8^2+5^2)=SQRT(64+25)=SQRT(89)
Answers are:
1) 53
2) 53 sqrt 2
3) 2 sqrt 53
4) sqrt 106 ...NONE OF THE ANSWERS ARE CORRECT..CHECK YOUR COORDINATES FROM GRAPH...APPLY THE ABOVE FORMULA...AB=SQRT{(DIFFERENCE IN X COORDINATES)^2+(DIFFERENCE IN Y COORDINATES)^2}